OpenStudy (anonymous):

Somebody please help? Find the inverse of f(x)=3x+2? And is that a function?!

OpenStudy (anonymous):

y= 3x +2 Switch X and Y x=3y +2 Divide 3 on both sides 1/3x= y+2/3 Isolate Y y= -1/3x +2/3 Btw: x/3 and 1/3x is the same thing

OpenStudy (anonymous):

yes

OpenStudy (cwrw238):

let y = 3x + 7 3x = y - 7 x = (y-7) / 3 f-1(x) = (x-7) / 3

OpenStudy (cwrw238):

oops - sorry - replace 7 by 2!!

OpenStudy (cwrw238):

yes - the inverse is a function

OpenStudy (anonymous):

seee christian im always right lol ;) ----Hmmm i sound like a stalker ^.^ -----LOl im not tho